OpenStudy (anonymous):

Please help with these 2 questions log x =log 2x^2-2 6^3x+1=215 Solve for x round to 2 decimal places

OpenStudy (agreene):

it's a bit hard to tell exactly what the questions are. Are they: \[\log (x)= \log(2x^2)-2\] and \[\large 6^{3x}+1=215\] ?